OpenStudy (godlovesme):

help please (: Are the two figures below similar? Explain why or why not using complete sentences. Provide evidence to support your claims.

OpenStudy (alexandervonhumboldt2):

i'm here

OpenStudy (godlovesme):

thanks :)

OpenStudy (alexandervonhumboldt2):

simmilar figures have congruent angles and proportional sides

OpenStudy (godlovesme):

i cant tell if the angles are congruent.... but idk how to prove if the sides are proportional :/

OpenStudy (alexandervonhumboldt2):

find lenghts of easch pair of corresponding sides and find the proportion between them. if this proportion will be same for all corresponding sides then these figures are similar.

OpenStudy (alexandervonhumboldt2):

use distance formula to find lenghts of sides

OpenStudy (godlovesme):

okay give me a sec i'll calculate the distance

OpenStudy (alexandervonhumboldt2):

hmm looks like if i remmebr corretly my brother my doing that test and it looks like they are not simmilar, i'm lazy calculation this all.

OpenStudy (godlovesme):

i will do the calculation :) just check mai answer please

OpenStudy (alexandervonhumboldt2):

ok

OpenStudy (alexandervonhumboldt2):

hmm well look like they are similar

OpenStudy (godlovesme):

CD= _/(3-(-1)^2 + (4-6)^2 CD= _/16 +4 CD= 2_/5

OpenStudy (godlovesme):

' _/ ' --> is the square root symbol xD

OpenStudy (anonymous):

my advice is: treat the sides like vectors and look how much they advance in vertical and horizontal directions. Example: BA = (2,-1), corresponding side of the smaller figure is IH=(1,-0.5) Do the same for other sides and look if there is any relation between corresponding sides components

OpenStudy (godlovesme):

im sorry but i dont get what u said .. :( sorry not good at this

OpenStudy (anonymous):

you know what vector is?

OpenStudy (anonymous):



OpenStudy (anonymous):

this vector that i just drowed, goes 3 right and 1 up. So we could write it like (3,1).

OpenStudy (anonymous):

 this second vector, goes 1 right an 1 down. We can write it as (1,-1). And so on

OpenStudy (anonymous):

Do the same for all the sides of your figures and compare numbers you get for corresponding sides. Look if you find any relation between them

OpenStudy (godlovesme):

i appreciate ur help @myko :) but i dont get this.. it's not you .. im kinda slow lol thanks thou(:

OpenStudy (anonymous):

kk